OpenStudy (anonymous):

7.38 times 10 ^2 ______________ 1.8 times 10^-3

OpenStudy (anonymous):

let's rewrite that to make it look nice :) \[\frac{7.38*10^2}{1.8*10^{-3}}\]

OpenStudy (anonymous):

You can bring the denominator to the top to make it positive.

OpenStudy (anonymous):

4.1x10^-1

OpenStudy (anonymous):

0.41

OpenStudy (anonymous):

That's not an option

OpenStudy (anonymous):

Don't give answers @JessicaPrajnadewi456

OpenStudy (anonymous):

Specially when they are wrong.

OpenStudy (anonymous):

\[\frac{7.38*10^2}{1.8*10^{-3}}\\ \frac{7.38*10^2*10^3}{1.8}\\ \frac{7.38*10^5}{1.8}\]

OpenStudy (anonymous):

Like i said bring it to the top then simplify

OpenStudy (anonymous):

This I would do it: calculate the different elements separately. 7.38/1.8=4.1. 10^2/10^-3 = 10^5. So your answer is 4.1*10^5

OpenStudy (anonymous):

Good job @BluFoot That's a good explanation